### v1

@@ Abstract
We introduce a new analytical framework for modelling degree sequences in individual communities of real-world networks, e.g., citations to papers in different fields. Our work is inspired by Price’s model and its recent generalisation called 3DSI (three dimensions of scientific impact), which assumes that citations are gained partly accidentally, and to some extent preferentially. Our generalisation is motivated by existing research indicating significant differences between how various scientific disciplines grow, namely, minding different growth ratios, average reference list lengths, and preferential citing tendencies. Extending the 3DSI model to heterogeneous networks with a community structure allows us to devise new analytical formulas for, e.g., citation number inequality and preferentiality measures. We show that the distribution of citations in a community tends to a Pareto type II distribution. We also present analytical formulas for estimating its parameters and Gini’s index. The new model is validated on real citation networks.
@@ 1 Introduction
In recent years, there has been a surge of interest in modelling real-life network data, especially since the rise of Graph Neural Networks Kipf and Welling (2017) and the emergence of open-access repositories of real networks, such as SNAP Leskovec and Krevl (2014), DBLP, or AMiner datasets. There is thus a growing need for theoretical network models, as they provide solid grounds for statistical reasoning and allow for simplifying assumptions about data.
Citation graphs are among the most prominently analysed and modelled networks: the publications represent vertices and bibliographic references are stored as directed edges. Such graphs help uncover patterns of knowledge dissemination, academic influence, and research trends over time Newman (2018) that are crucial topics in bibliometrics and science of science Milojević (2025). We can model citation networks with the help of theoretical random graph models, such as the renowned Barabási–Albert (2002) model, the Stochastic Block Model and its degree-corrected version Lee and Wilkinson (2019), or generatively, with tools such as GSCALER Zhang and Tay (2016) or the LFR Benchmark Lancichinetti et al. (2008).
@@
In our work, we are specifically interested in modelling real networks as iteratively growing random graphs with a community (cluster, subgroup) structure. The communities play a significant role in many types of real graphs, such as biological, urban Fortunato (2010), and social networks Chen et al. (2014), but their role in modern modelling of citation networks, where there is a division to scientific fields and knowledge domains, is often underplayed. Also, the problem of community detection has for a long time been prevalent in the area of network science Fortunato (2010). While the problem of finding communities is long-standing and we already have access to highly effective methods of finding communities, such as the Louvain Blondel et al. (2008) and Leiden Traag et al. (2019) algorithms, they cannot directly help us distinguish which network parameters govern the emergence of community structure and how the communities change over time. As such, we wish to propose a new network model with a community structure with a focus on the flexibility and interpretability of the underlying parameters. Continuing, we shortly describe the core areas in literature related to our model: the foundational models of citation network science, the role of communities in networks, and key modern approaches to citation network modelling.
Foundational models of network growth.
These models were the first and the most ground-breaking and sought to explain the scale-free nature of citation distributions. Here we should name, among others, Price’s model of cumulative advantage Price (1965) which was one of the first to introduce preferential attachment and describing the “rich-get-richer” phenomenon, i.e., that the probability that an existing paper receives a citation is directly proportional to the number of citation it already has. Further, the Barabási–Albert model of preferential attachment Albert and Barabási (2002) generalised the Price model and proved its further mathematical properties. Also, it highlighted the importance of the power-law distributions in many real networks, not only bibliographic ones. We should mention the Bianconi–Barabási fitness model Bianconi and Barabási (2001) too: it assumes that each node is also associated with an intrinsic “fitness” (e.g., quality of the newly published paper) that impacts the rate at which it gains citations. In particular, the fitness model helps understand the discovery of “sleeping beauties”, i.e., publications that are being dismissed for a long time only to then acquire a large number of citations over a short period.
Importance of communities.
Very quickly it became obvious that communities play an important role in the growth and modelling of many heterogeneous networks Fortunato (2010). One of the first models to deal with such networks is the Stochastic Block Model Holland et al. (1983); it is a generative statistical model for networks with community structure. It assumes that all nodes in a network belong to one of $k$ latent (unobserved) groups or “blocks”. The probability of an edge between two nodes depends only on the blocks to which they belong. Nowadays, SBM has been generalised in many directions, e.g., as parts of the Exponential Random Graph Model families Fronczak et al. (2013) (see below). Another group of models that take into account the role of the communities are topic models, such as the Latent Dirichlet Allocation, which assumes that papers are a mixture of topics and each topic is a probability distribution over words Blei et al. (2003). Using this model, we can represent each paper as a vector of topic proportions and then predict a citation link between any pair of papers based on the similarity of topic distributions. For example, the Relational Topic Model (RTM) Chang and Blei (2009) directly combines LDA and link prediction models. Also, Drobyshevskiy and Turdakov (2019) published a comprehensive survey of random graph models in which they include eight random graph models capable of generating networks with a community structure. For the survey of models not focused on citation networks, we direct the reader to their work. We should note that despite the extensiveness of this survey, neither of the considered models allowed for community structure under the preferential attachment rule. This, indirectly, emphasises the novelty of our model.
@@
Many modern approaches generalise, in some way, the previously mentioned core models or focus on data-driven generation of new graphs. Here we can mention Graph Neural Networks Kipf and Welling (2017) or Graph Variational Autoencoders Li et al. (2020). However, these models often lack statistical precision and elegance and are based on heuristic assumptions. Another approach gaining popularity is comprised of the aforementioned Exponential Random Graph Models (ERGMs), which, using a statistical framework, define a probability distribution over the space of all possible graphs of a certain size. While highly efficient in testing hypotheses on the networks, they still require computational-heavy fitting to real data. On the other end of the spectrum are the models such as the 3DSI (three dimensions of scientific impact) model Siudem et al. (2020), which we describe in more detail below. It generalises the standard procedure of iterative preferential gains of the Barabási–Albert and Price models by adding an additional component of accidental attachments. It is easy to fit to the data and offer valuable analytical insight into the data, but in its standard formulation it does not allow communities in the data.
@@
We prove some theoretical results regarding, amongst others, the Gini index of the node degree distribution, as well as some asymptotic properties (Section 2.4). In Section 3, we demonstrate the new model’s usefulness in describing real-world citation networks. Section 4 concludes the paper.
@@ 2.1 Homogeneous Price–Pareto model
Here are the basic assumptions behind the standard 3DSI citation network growth presented in Siudem et al. (2020). It was based on the Price model and further extended in Siudem et al. (2022).
@@ 2.2 Introducing community structure
where $\operatorname*{Acc}_{i}^{(t)}(\ell)$ denotes the expected accidental income of the node $\ell$ at local turn $t$ assuming that $\ell$ belongs to the community $i$. To calculate $\operatorname*{Acc}_{i}^{(t)}(\ell)$, we make a few observations:
@@
    Since the network is growing, we consider the accidental income for the node $\ell$ at each of the global turns $X+s$ for some $s\geq 0$. The variable $s$ increases as long as we do not draw a new vertex from the community $i$. So $(1-p_{i})^{s-1}$ is the probability that the number of global steps between the local time t1t-1 and $t$ is equal to $s$.
@@
    Since the community probabilities are independent of $X$ and $t$, we can write that the expected accidental income of the vertex $\ell$ in the global turn $X+s$ is equal to:
@@
    What follows is that the total expected accidental income of the node $\ell$ in the local turn t1t-1 is equal to:
@@
    $=\langle a\rangle\int_{0}^{1}\sum_{s=0}^{\infty}(1-p_{i})^{s}u^{s}\mathbb{E}_{X}\left[u^{X-1}\right]du$
@@
    $\left[z=(1-p_{i})u,dz=(1-p_{i})du\right]$ $=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\int_{0}^{1-p_{i}}\frac{z^{t-2}}{(1-z)^{t}}dz$
    $\left[y=\frac{z}{1-z},dy=\frac{dz}{(1-z)^{2}}\right]$ $=\langle a\rangle\int_{0}^{\frac{1-p_{i}}{p_{i}}}y^{t-2}dy=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\frac{(1-p_{i})^{t-1}}{p_{i}^{t-1}}\frac{1}{t-1}$
@@
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\frac{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle+\rho_{i}m_{i}}{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle}=$
@@
Figure 1: Example model-predicted node in-degree sequences for communities in a graph with $N=30000$ nodes and three communities
@@
Figure 1 presents the node in-degrees output by our model for a graph with three communities on $30000$ vertices. We observe that modifying different parameters has impact on different aspects of the final sequence: higher $m_{i}$ mostly impacts the overall slope of the sequence, $\rho_{i}$ impacts the curvature of the degree sequence the most, while $p_{i}$ changes the amplitude of the sequence as it directly represents the number of vertices in the community.
@@ 2.3 Gini index
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\sum_{\ell=1}^{t}(t-2\ell+1)\left[\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}-1\right]=$
@@
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[\frac{(t+1)(2-\nu_{i})}{(1-\nu_{i})(2-\nu_{i})}+\frac{2\nu_{i}t-2t-2}{(1-\nu_{i})(2-\nu_{i})}\right]=$
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\frac{2t-\nu_{i}t+2-\nu_{i}+2\nu_{i}t-2t-2}{(1-\nu_{i})(2-\nu_{i})}=$
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\frac{\nu_{i}(t-1)}{(1-\nu_{i})(2-\nu_{i})}=\frac{\langle a\rangle}{\Sigma_{i}^{(t)}}\frac{t-\nu_{i}}{(1-\nu_{i})(2-\nu_{i})}=$
$=\frac{t-\nu_{i}}{t-1}\frac{1}{2-\nu_{i}},$ (8)
where the last two equalities follow by applying Equation (5) and Equation (6). Of course, for large $t$, $\mathcal{G}_{i}^{(t)}\approx\frac{1}{2-\nu_{i}}$.
As with the global degree sequence, finding the Gini coefficient of the whole sequene is difficult due to the mixture nature of the modelled sample. We provide clean asymptotic results though, which we derive in the next section.
@@ 2.4 Asymptotic results: Price meets Pareto
We can consider the theoretical formulas asymptotically for $t\rightarrow\infty$, as both network sizes and community sizes are often sufficiently big. Following Siudem et al. (2022), we consider the inverse of the survival function:
@@
$f_{i}(x)=\frac{dF_{i}(x)}{dx}=\frac{1}{\langle a\rangle}\left(1+\frac{\nu_{i}}{\langle a\rangle}x\right)^{-1-\frac{1}{\nu_{i}}}.$
@@
Numerical tests conducted on a battery of samples with various sets of parameters and $10000\leq N\leq 20000$ indicate that the difference between the Gini coefficient value calculated with the exact formula and with the approximation almost never exceeds $0.02$ and that on average the difference is very close to 0. While Siudem et al. (2022) also presents clean derivation of the order statistics for the standard 3DSI degree sequence, we have found that introducing the community mixture makes the resulting formulas unsimplifiable (though they can be represented using, i.a., the Lauriciella generalised hypergeometric functions). As such, we omit this derivation here.
@@ 3.1 Datasets and data curation
We fit our model to two datasets: the CORA database of scientific papers classified into seven disciplines Sen et al. (2008), and the DBLP v14 citation network Tang et al. (2008). We use the fourteenth edition of the dataset as it is the latest one to contain information about fields of research, which we use as a source for the ground-truth communities. While the standard DBLP v14 dataset is a citation network of research articles, we aggregated it to obtain a citation network where the nodes represent authors, and the unweighted edges between authors represent citations. The ground-truth community of each author was selected as the community with the highest total discipline weight from all their papers. An edge between two authors is present if one cites the other at least once. We also remove self-citations in accordance with the model’s assumptions. With this procedure, we try to reduce the noise coming from the automated discipline assignment to each paper. We also restrict the data from the DBLP datasets to the network induced by the eight largest communities, which we believe is sufficient for the presentation of the results while avoiding unnecessary clutter. We performed no additional data transformation on the CORA dataset. Table 1 contains a general description of the two datasets.
@@
CORA 2708 5429 7 2.005
DBLP V14 Authors 481387 58544370 8 121.616
Figure 2: Observed and modelled in-degree densities in the eight largest clusters of the DBLP authors dataset. The bar plot with orange dots presents the observed density, while the lines presents the estimated densities from our model corresponding to the Gini-based $\rho_{i}$ estimation. The last plot shows the full in-degree sequences and the estimations.
Figure 3: Observed and modelled in-degree densities in the seven clusters of the CORA dataset. The bar plot with orange dots presents the observed density, while the lines presents the estimated densities from our model corresponding to the Gini-based $\rho_{i}$ estimation. The last plot shows the full in-degree sequences and the estimations.

### v2

@@ Abstract
We introduce a new analytical framework for modelling degree sequences in individual communities of real-world networks, e.g., citations to papers in different fields. Our work is inspired by a recent modification of the Price’s model, which assumes that citations are gained partly accidentally, and to some extent preferentially. Our work addresses the need to represent the heterogeneity of various scientific domains, as standard homogeneous models fail to capture the distinct growth ratios and citing cultures of different fields. Extending the model to networks with a community structure allows us to devise the analytical formulae for, amongst others, citation counts in each cluster and their inequality as described by the Gini index. We also show that a citation count distribution in each community tends to a Pareto type II distribution. Thanks to the derived model parameter estimators, the new model can be fitted to real citation and similar networks.
@@ 1 Introduction
In recent years, there has been a surge of interest in analysing real-life network data, stimulated by the development of neural architectures for graph processing, such as Graph Neural Networks Scarselli et al. (2009) and Graph Convolutional Networks Kipf and Welling (2017), as well as the emergence of open-access repositories of real networks, such as SNAP Leskovec and Krevl (2014), DBLP, or AMiner datasets. There is thus a growing need for theoretical network models, as they provide solid grounds for statistical reasoning and allow for simplifying assumptions about data. Citation graphs are among the most prominently analysed and modelled networks: the publications represent vertices and bibliographic references are stored as directed edges. Such graphs help uncover patterns of knowledge dissemination, academic influence, and research trends over time Newman (2018) that are crucial topics in bibliometrics and science of science Milojević (2025).
@@
In our work, we are specifically interested in modelling real networks as iteratively growing random graphs with a community (cluster, subgroup) structure. The communities play a significant role in many types of real graphs, such as biological, urban Fortunato (2010), and social networks Chen et al. (2014), but their role in modern modelling of citation networks, where there is a division to scientific fields and knowledge domains, is often underplayed. Also, the problem of community detection has for a long time been prevalent in the area of network science Fortunato (2010). While the problem of finding communities is long-standing and we already have access to highly effective methods of finding communities, such as the Louvain Blondel et al. (2008) and Leiden Traag et al. (2019) algorithms, they cannot directly help us distinguish which network parameters govern the emergence of community structure and how the communities change over time.
As such, we wish to propose a new network model featuring a community structure with a focus on the flexibility and interpretability of the underlying parameters. Our model can be applied to any network with either exogenous or algorithmically found communities, as long as they promote associative connections; that is, we require that citations are more likely to happen between two works from the same community. Also, while our model is derived on the ground of paper-to-paper citation networks, we also consider its applications to, e.g., author-to-author networks. To formulate our model concisely and keep the number of parameters reasonable, we formulate our proposition as a minimal analytical extension of an existing 3DSI model Siudem et al. (2020) which we describe below. Our generalisation is motivated by the necessity to model the heterogeneity inherent in scientific landscapes. Existing research indicates significant differences between how various scientific disciplines (represented here as communities or clusters) grow, characterised by distinct growth ratios, average reference list lengths, and preferential citing tendencies Cascarina (2023); Harzing (2010); Simkin and Roychowdhury (2007). A single global parameter set cannot capture these local dynamics; for instance, a citation in a fast-moving field like Artificial Intelligence has different dynamical characteristics than one in Classical Logic.
Continuing, we shortly describe the core areas in literature related to our model: the foundational models of citation network science, the role of communities in networks, and key modern approaches to citation network modelling.
Foundational and citation-specific growth models.
The earliest attempts to explain the scale-free nature of citation distributions focused on the “rich-get-richer” phenomenon. Price’s model of cumulative advantage Price (1965) was among the first to formalise preferential attachment, a concept later generalised by the Barabási–Albert model Albert and Barabási (2002), which highlighted the ubiquity of power-law distributions in complex networks. However, simple preferential attachment inevitably leads to a strong bias towards older nodes. To address this, fitness-based models were introduced, such as the Bianconi–Barabási model Bianconi and Barabási (2001), which assigns an intrinsic “fitness” to nodes to explain how younger papers can overtake older ones (e.g., “sleeping beauties”). More recently, specific attention has been paid to the temporal dynamics of citations. Models proposed by Medo et al. (2011) and Golosovsky (2017) explicitly incorporate mechanisms for node relevance decay, demonstrating that the probability of receiving a citation must diminish over time to match empirical observations. This aligns with bibliometric studies on “attention decay” Parolo et al. (2015) and the long-term impact of scientific works Wang et al. (2013), which show that citation rates follow a rise-and-fall pattern not fully captured by standard scale-free models.
Modelling communities and structure.
While growth models focus on temporal dynamics, a parallel stream of research focuses on the modular structure of networks. The Stochastic Block Model (SBM) Holland et al. (1983) serves as the generative baseline for networks with latent group structures. Unlike growth models, SBMs are typically static, but they offer rigorous connections between model parameters and structural properties; notably, Decelle et al. (2011) identified precise phase transitions in SBMs that dictate when community structure becomes theoretically detectable. Other approaches for generating clustered networks include the LFR Benchmark Lancichinetti et al. (2008) and GSCALER Zhang and Tay (2016). While these tools are invaluable for benchmarking community detection algorithms, they are designed primarily to reproduce static structural features (e.g., degree distributions and community sizes) rather than to model the process of network growth via citation dynamics. Topic models, such as Latent Dirichlet Allocation (LDA) Blei et al. (2003) and the Relational Topic Model (RTM) Chang and Blei (2009), bridge this gap by representing papers as vectors of topic proportions and predicting links based on the similarity of these distributions. However, these approaches often lack the analytical simplicity of pure random graph models. Notably, Drobyshevskiy and Turdakov (2019) published a comprehensive survey including eight random graph models capable of generating networks with a community structure, yet they found that none allowed for community structure under the preferential attachment rule. This highlights a critical gap in analytical models that simultaneously capture the temporal growth of citation networks and their heterogeneous community structure.
@@
Many modern approaches generalise the previously mentioned core models or focus on data-driven generation of new graphs. Here we can mention Graph Neural Networks Kipf and Welling (2017) or Graph Variational Autoencoders Li et al. (2020). However, these models often lack statistical precision and elegance and are based on heuristic assumptions. Another approach gaining popularity is comprised of Exponential Random Graph Models (ERGMs), which can be viewed as generalisations of SBMs Fronczak et al. (2013). Using a statistical framework, they define a probability distribution over the space of all possible graphs of a certain size. While highly efficient in testing hypotheses on the networks, they still require computational-heavy fitting to real data. On the other end of the spectrum are models such as the 3DSI (three dimensions of scientific impact) model Siudem et al. (2020); see Section 2.1. It generalises the standard procedure of iterative preferential gains of the Barabási–Albert and Price models by adding an additional component of accidental attachments. The name comes from the model’s ability to describe citation records using only three parameters: productivity, total impact, and how lucky an author has been so far. It is easy to fit to the data and offer valuable analytical insight into the data, but in its standard formulation it does not allow communities in the data.
@@
We prove some theoretical results regarding, amongst others, the Gini index of the node degree distribution, as well as some asymptotic properties, such as the limiting distribution’s being Paretian (Section 2.4). In Section 3, we demonstrate the new model’s usefulness in describing real-world citation networks. Section 4 concludes the paper.
@@ 2.1 Homogeneous Price–Pareto model
Here are the basic assumptions behind the standard 3DSI citation network growth presented in Siudem et al. (2020).
@@
Conceptually, 3DSI is similar to the Price (1965) power-law-type model. However, in the latter, the preferential citation gain of node $\ell$ is proportional to $d_{i}^{(t)}(\ell)+C$ for some constant $C$. On the other hand, the key assumption behind the 3DSI model is that in each turn, first the accidental citations are added, and only then the preferential ones. Thanks to this simplification, the constant $C$ is not necessary, since after the accidental distribution of citations the expected in-degree of each article is non-zero. From an analytical perspective, this is a notable improvement over Price’s approach, because the constant $C$ was generally difficult to both estimate and interpret Milojević (2020). Moreover, as we will mention below, asymptotically, 3DSI enjoys convergence to a Pareto type II distribution. To account for its standing on the shoulders of giants, we will also be referring to 3DSI as the Price–Pareto model.
@@ 2.2 Introducing community structure
Notation.
In the following derivations, let $t$ denote the discrete local time step in the $i$-th community ($t_{i}$ suppressing the subscript), corresponding directly to the number of vertices currently assigned to that community. Let $d_{i}^{(t)}(\ell)$ denote the in-degree of a vertex $\ell$ belonging to community $i$ at local time $t$. Finally, let $\operatorname*{Acc}_{i}^{(t)}(\ell)$ denote the expected accidental gain accumulated by vertex $\ell$ in community $i$ during the interval between local time steps t1t-1 and $t$. In the derivations, we frequently make use of the fact that the vertex identifier corresponds to its entry time.
@@
To calculate $\operatorname*{Acc}_{i}^{(t)}(\ell)$, we make a few observations:
@@
    Since the network is growing, we consider the accidental gain for the node $\ell$ at each of the global turns $X+s$ for some $s\geq 0$. The variable $s$ increases as long as we do not draw a new vertex from the community $i$. So $(1-p_{i})^{s-1}$ is the probability that the number of global steps between the local time t1t-1 and $t$ is equal to at least $s$.
@@
    Since the community probabilities are independent of $X$ and $t$, we can write that the expected accidental gain of the vertex $\ell$ in the global turn $X+s$ is equal to:
@@
    What follows is that the total expected accidental gain of the node $\ell$ in the local turn t1t-1 is equal to:
@@
    $=\langle a\rangle\int_{0}^{1}\sum_{s=0}^{\infty}(1-p_{i})^{s}u^{s}\mathbb{E}_{X}\left[u^{X-1}\right]du=$
@@
    $=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\int_{0}^{1-p_{i}}\frac{z^{t-2}}{(1-z)^{t}}dz=$
    $=\langle a\rangle\int_{0}^{\frac{1-p_{i}}{p_{i}}}y^{t-2}dy=$
@@
Figure 1: Example model-predicted node in-degree sequences for communities in a graph with N=30, [^0] N=30{,}000 nodes and three communities
@@
Figure 1 presents the node in-degrees output by our model for a graph with three communities on 30,00030{,}000 vertices. We observe that modifying different parameters has impact on different aspects of the final sequence: higher $m_{i}$ mostly impacts the overall slope of the sequence, $\rho_{i}$ impacts the curvature of the degree sequence the most, while $p_{i}$ changes the amplitude of the sequence as it directly represents the number of vertices in the community.
@@ 2.3 Gini index
$=\frac{\langle a\rangle}{\Sigma_{i}^{(t)}}\frac{t-\nu_{i}}{(1-\nu_{i})(2-\nu_{i})}=\frac{t-\nu_{i}}{t-1}\frac{1}{2-\nu_{i}},$ (8)
where the last equality follows by applying Equation (5). Of course, for large $t$, $\mathcal{G}_{i}^{(t)}\approx\frac{1}{2-\nu_{i}}$.
As with the global degree sequence, finding the Gini coefficient of the whole sequence is difficult due to the mixture nature of the modelled sample. We provide clean asymptotic results though, which we derive in the next section.
@@ 2.4 Asymptotic results: Price meets Pareto
We can consider the theoretical formulae asymptotically for $t\rightarrow\infty$, as both network sizes and community sizes are often sufficiently big. Following Siudem et al. (2022), we consider the inverse of the survival function:
@@
$f_{i}(x)=\frac{dF_{i}(x)}{dx}=\frac{1}{\langle a\rangle}\left(1+\frac{\nu_{i}}{\langle a\rangle}x\right)^{-1-\frac{1}{\nu_{i}}},$
@@
Numerical tests conducted on a battery of samples with various sets of parameters and $10{,}000\leq N\leq 20{,}000$ indicate that the difference between the Gini coefficient value calculated with the exact formula and with the approximation almost never exceeds $0.02$ and that on average the difference is very close to 0. While Siudem et al. (2022) also presents clean derivation of the order statistics for the standard 3DSI degree sequence, we have found that introducing the community mixture makes the resulting formulae unsimplifiable (though they can be represented using, i.a., the Lauriciella generalised hypergeometric functions). As such, we omit this derivation here.
@@ 3.1 Datasets and data curation
We fit our model to two well-established datasets. First, we utilise the Cora dataset Sen et al. (2008), a benchmark in machine learning research consisting of 2,708 scientific papers classified into seven distinct sub-disciplines (e.g., Neural Networks, Genetic Algorithms). This dataset represents a standard paper-to-paper citation network where edges are directed citations. It is widely used as a benchmark dataset in tasks such as node classification.
Second, we analyse the DBLP v14 dataset Tang et al. (2008), a comprehensive bibliographic database of computer science publications. We use the fourteenth edition as it contains field-of-study information, which we use as the source for ground-truth communities. To study an author-to-author topology, we projected the paper-to-paper graph into a simple directed graph where nodes represent paper authors. Specifically, a directed edge is created from author $u$ to author $v$ if author $u$ has written a paper that cites any paper written by author $v$. To adhere to our model’s assumptions (simple directed graph), we flattened the resulting multigraph by treating multiple citations between the same pair of authors as a single unweighted edge and removed self-loops (self-citations). The ground-truth community of each author was determined by weighted majority vote: the author is assigned to the community corresponding to the discipline with the highest total weight over all the papers of this author, where the weight was extracted from the field-of-study information in the data. The motivation for the transformation we performed on DBLP is two-fold: first, it allows us to reduce the noise in ground-truth community labels that were assigned automatically; second, it allows us to test our model on an author-to-author network instead of a standard paper-to-paper network.
@@
CORA 2,708 5,429 7 2.005
DBLP V14 Authors 481,387 58,544,370 8 121.616
Figure 2: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the DBLP authors dataset. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.
Figure 3: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the Cora dataset. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.